OpenStudy (anonymous):

A university bookstore recently sold a wirebound graph-paper notebook for $2.55, and a college-ruled notebook for $3.22. At the start of spring semester, a combination of 50 of these notebooks were sold for a total of $157.65. How many of each type were sold?

OpenStudy (anonymous):

this is a system, you have to create two equations

OpenStudy (anonymous):

i would begin by lableing x as wire bound graph notebook and y as college ruled notebook

OpenStudy (anonymous):

now that we have our our two variables, we can begin with the easy equation: the combination of notebooks being 50, this means: x+y=50

OpenStudy (anonymous):

now our second equation comes from the price of each notebook, x (the graph paper) cost 2.55 dollars. The y (wide ruled paper) cost 3.22 dollars. and the total cost of the notebooks is 157.65 dollars. Our mathematical equation for this is: 2.65x + 3.22y = 157.65

OpenStudy (anonymous):

now we have our system: x+y=50 2.65x + 3.22y = 157.65

OpenStudy (anonymous):

now simply solve this system